# Q-BERT4Rec: Quantized Semantic-ID Representation Learning for Multimodal Recommendation

### v1

@@ 1. INTRODUCTION
Recommend systems have become indispensable across e-commerce, digital advertising, and content streaming, where they surface relevant items from massive catalogs and drive key business metrics such as CTR, retention, and revenue(chen2020survey). Among various paradigms, sequential recommendation focuses on modeling users’ historical interactions to predict their next actions, thereby capturing evolving preferences and short-term intents. Early neural methods such as GRU4Rec(jannach2017recurrent) leveraged recurrent architectures to encode temporal dynamics, while self-attention models like SASRec(kang2018self) improved long-range dependency modeling and parallelism. Building on the Transformer, BERT4Rec(sun2019bert4rec) introduced bidirectional contextualization for sequences and set a strong foundation for many subsequent advances. These developments establish the Transformer as a de facto backbone for sequence modeling in recommendation.
@@
To enhance representation quality, recent work focuses on exploring multimodal signals—text (titles, descriptions), images, and structured attributes—to enrich item semantics beyond bare IDs (e.g., VBPR, MMGCN, DualGNN)(he2016vbpr; wei2019mmgcn; wang2021dualgnn). In parallel, generative and prompt-based recommenders reframe recommendation as a language or sequence generation problem(deldjoo2024review), unifying diverse tasks through instruction tuning or tokenization (e.g., P5, M6-Rec, TIGER, VIP5)(geng2022recommendation; cui2022m6; rajput2023recommender; geng2023vip5). More recent work such as MQL4GRec(zhai2025multimodal) and DiffMM(jiang2024diffmm)further explored quantized or generative modeling to bridge continuous multimodal features and discrete semantics. MQL4GRec employs Residual Vector Quantization (RQ-VAE) (zhai2025multimodal) to convert multimodal representations into discrete semantic tokens. However, it performs quantization independently for each modality, which results in inconsistent codebook distributions across modalities. This misalignment weakens the shared semantic space and hinders the integration of multimodal information in sequential modeling. DiffMM, while leveraging diffusion-based reconstruction(ho2020denoising) for generation, primarily focuses on denoising processes rather than fine-grained multimodal fusion. These examples reveal a fundamental challenge: existing models either perform static multimodal fusion without adaptivity, or decouple quantization from sequence modeling, thereby limiting both interpretability and generalization.
@@
The development of Residual Vector Quantization (RQ-VAE) techniques(oord2017vqvae; zeghidour2021soundstream) has further inspired us to address these limitations. Similarly to how pretraining–finetuning(devlin2019bert) and prompt-tuning(brown2020gpt3) revolutionized natural language generation by introducing a shared vocabulary with transferable semantics, we envision a comparable paradigm shift for recommendation. Instead of representing items by arbitrary, meaningless IDs, we propose to replace them with semantic IDs—compact sequences of quantized semantic tokens that encapsulate multimodal meaning. These semantic IDs act as both identifiers and language units, bridging the gap between discrete recommendation tokens and rich multimodal semantics. In this view, each token from the quantized codebook serves as a semantic atom, much like words in a vocabulary that encode transferable priors across domains. Consequently, recommendation can be reformulated as a language of items, where interactions are modeled as sentences composed of semantic tokens. This new representation not only enables cross-domain knowledge transfer and interpretability, but also maintains efficiency and scalability by being more compact than original modalities such as text or images. As illustrated in Figure 1, the proposed semantic-ID paradigm provides a unified, token-level bridge connecting multimodal content, quantized representations, and sequential reasoning.
@@
To this paper, we propose a novel Quantized BERT-style Multimodal Sequential Recommendation framework, named Q-Bert4Rec. Specifically, our method is motivated by the observation that traditional sequential recommenders represent items as discrete and meaningless IDs, which lack semantic understanding and hinder generalization across domains. To bridge this gap, Q-Bert4Rec introduces a unified three-stage framework that learns semantic IDs composed of quantized multimodal tokens, enabling both semantic richness and discrete modeling. In the first stage, we perform Cross-modal Semantic Injection, where a dynamic Transformer fuses textual, visual, and structural features into randomly initialized ID embeddings. A learnable gating mechanism adaptively controls the fusion depth, ensuring content-aware semantic enrichment for different items.In the second stage, we introduce Semantic Quantization using a Residual Vector Quantization module (RQ-VAE)(zeghidour2021soundstream). The fused representations are discretized into compact, interpretable semantic tokens, which serve as the new identifiers—semantic IDs—to replace original item IDs. These quantized tokens form a unified vocabulary that encodes multimodal meaning in a transferable and efficient way. In the third stage, we design a Multi-mask Pretraining and Fine-tuning strategy to fully exploit sequential dependencies. By combining span, tail, and multi-region masking, the model captures both local continuity and long-range correlations in user behavior sequences. This stage enhances the model’s ability to reason over diverse sequential patterns and improves its adaptability to downstream recommendation tasks. We evaluate Q-Bert4Rec on multiple public Amazon benchmarks, and the results show significant improvements compared to strong baselines. The findings demonstrate that unifying semantic fusion and quantized modeling effectively bridges the gap between continuous multimodal representations and discrete recommendation reasoning.
@@ 2. RELATED WORK
Sequential recommendation. Sequential models aims to capture users’ dynamic interaction histories and predict their next potential actions. Early neural architectures, such as GRU4Rec (jannach2017recurrent) and Caser (tang2018personalized), leveraged recurrent and convolutional structures to capture temporal patterns. Later, the emergence of self-attention mechanisms, particularly SASRec (kang2018self), greatly improved long-range dependency modeling and parallel training efficiency. Building upon this, BERT4Rec (sun2019bert4rec) introduced a bidirectional Transformer encoder and masked item prediction, setting a strong foundation for subsequent Transformer-based recommenders.Recent works have explored different enhancements on this backbone. For instance, CL4SRec (xie2022contrastive) and DuoRec (qiu2022contrastive) introduced contrastive learning objectives to improve representation robustness, while FDSA (zhang2019feature) incorporated feature-level attention to exploit auxiliary attributes. Despite their success, these approaches still represent items using discrete, non-semantic IDs, which limits their transferability to unseen domains or new items. To address this, a new research trend focuses on generative recommendation, which reformulates recommendation as a language modeling or sequence generation problem. Representative works include P5 (geng2022recommendation), M6-Rec (cui2022m6), and TIGER (rajput2023recommender). These models leverage pre-trained language models (PLMs) and design prompt-based tasks to unify multiple recommendation objectives. P5 and M6-Rec transform recommendation into text generation through task prompts, while P5-ID (hua2023index) explore how the design of item identifiers affects model transferability. ColaRec (wang2024enhanced) constructs generative item IDs by encoding collaborative signals, whereas TIGER(rajput2023recommender) is the first to introduce RQ-VAE(zeghidour2021soundstream) for constructing discrete item IDs via vector quantization.
@@
Multimodal recommendation. Multi-modal models exploit side information—such as item text, images, and metadata—to provide richer representations of user–item interactions. Early works, such as VBPR (he2016vbpr), enhanced matrix factorization with visual features extracted from product images. Subsequent research introduced graph-based models to capture modality relationships and high-order dependencies. For example, MMGCN (wei2019mmgcn) used graph neural networks to propagate multimodal signals; LATTICE (zhang2021mining) built modality-specific item-item graphs for structural fusion; and DualGNN (wang2021dualgnn) modeled cross-modal attentions between user preferences and item content. Similarly, MVGAE (yi2021multi) employed modality-specific variational autoencoders to integrate multimodal embeddings in a unified latent space. With the rapid advancement of large pre-trained models, multimodal recommendation has shifted toward leveraging foundation encoders for representation learning. VIP5 (geng2023vip5) extends the P5 framework by incorporating CLIP-based image encoders to align visual and textual modalities within a generative prompting framework. MMGRec (liu2024mmgrec) further utilizes a Graph RQ-VAE to construct item IDs from multimodal and collaborative signals, unifying item quantization and relational reasoning. IISAN (fu2024iisan) proposes a decoupled PEFT architecture for lightweight multimodal adaptation, demonstrating the scalability of pretrained multimodal encoders.
@@ 3.1. Problem Formulation
This bridges continuous multimodal semantics and discrete token-level modeling, enabling interpretability and efficient knowledge transfer.
@@ 3.3. Dynamic Cross-Modal Semantic Injection
(5) $\mathcal{L}_{\text{align}}=\mathcal{L}_{it}+\mathcal{L}_{iv}+\mathcal{L}_{tv}+\lambda\,\|\mathbf{h}_{i}-\text{sg}(\mathbf{q}_{i})\|_{2}^{2},$
where $\mathcal{L}_{it}$, $\mathcal{L}_{iv}$, and $\mathcal{L}_{tv}$ are symmetric InfoNCE losses between the fused, text, and image embeddings, and $\text{sg}(\cdot)$ denotes the stop-gradient operator. The first three terms encourage cross-modal alignment, while the consistency term preserves the ID semantics and stabilizes training.
@@ 3.4. Semantic Quantization
Handling collisions. In practice, multiple items may be quantized to identical token sequences, leading to code collisions, where different items have the same token sequences. To mitigate this, we adopt the reallocation strategy proposed in MQL4GRec (zhai2025multimodal). For $N$ colliding items, we first compute the distance tensor $\mathbf{D}\in\mathbb{R}^{N\times L\times K}$ between residual vectors and codewords in each level:
@@ 4.1. Experimental Setting
Dataset. We conduct experiments on the widely used Amazon Product Reviews dataset(ni2019justifying), which contains user–item interaction histories of different domains collected from Amazon over several years. Each domain represents a distinct product category and includes user reviews, ratings, and timestamps, making it suitable for evaluating cross-domain sequential recommendation.Statistics of these datasets are shown in Table  1.
@@
Evaluation Metrics. We evaluate model performance using two widely adopted ranking metrics: Hit Ratio (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K) with K = 1, 5, 10. Following previous work(kang2018self), We use the leave-one-out for evaluation, where the most recent interaction is held out for testing and the second most recent for validation.
Implementation Details. We implement our model using PyTorch. For the dynamic cross-modal semantic injection module, we employ the pretrained LLaMA model to encode textual representations, and adopt the image encoder branch from CLIP (radford2021learning), which uses ViT (dosovitskiy2020image) as its backbone, to extract visual features of items. The Transformer within this module is configured with a maximum depth of 5 layers.
@@
Following prior work (sun2019bert4rec), during pretraining, we apply three masking strategies with probabilities of 0.3 (span mask), 0.15 (multi-position mask), and 0.1 (tail mask). The model is trained with a batch size of 512 and a learning rate of 0.0005. The Transformer encoder consists of 4 layers with 4 attention heads. For fine-tuning, we use only the traditional mask strategy with a ratio of 0.3, a batch size of 256, and a learning rate of 0.0001. To prevent overfitting, we apply dropout with a rate of 0.2 to all Transformer layers.
@@ 4.2. Performance Comparison (RQ1)
Baselines. In this section, we compare our proposed approach with a wide range of sequential recommendation baselines, which are briefly described in Appendix A. The compared methods include GRU4Rec(jannach2017recurrent), BERT4Rec (sun2019bert4rec), SASRec(kang2018self), FDSA(zhang2019feature), S³-Rec(zhou2020s3), VQ-Rec(hou2023learning), MISSRec(wang2023missrec), P5-CID(hua2023index), VIP5(geng2023vip5), TIGER(rajput2023recommender), and MQL4Grec(zhai2025multimodal). The experimental results are reported in Table 2.
@@
These consistent improvements validate that dynamic cross-modal semantic injection effectively captures richer multimodal dependencies, while the residual quantization module enhances semantic compactness and generalization across domains. Furthermore, the multi-mask pretraining strategy strengthens the model’s ability to understand diverse user interaction patterns, leading to better next-item prediction accuracy. Overall,in answer to RQ1, Q-BERT4Rec achieves the best performance in 13 out of 15 evaluation metrics, demonstrating strong generalization, interpretability, and robustness across heterogeneous recommendation domains.
@@ Appendix A Baselines
  • GRU4Rec (jannach2017recurrent) employs gated recurrent units to model sequential dependencies in user behaviors. Although effective for short sequences, it struggles to capture long-range dependencies and contextual semantics.
@@
  • BERT4Rec (sun2019bert4rec) applies bidirectional Transformers with the Masked Language Modeling (MLM) objective for next-item prediction. However, it models discrete item IDs without semantic meaning.
  • SASRec (kang2018self) uses uni-directional self-attention for autoregressive sequence modeling. It effectively captures local dependencies but ignores item-level multimodal semantics.
  • FDSA (zhang2019feature) integrates user and item attributes into the attention mechanism to enhance personalization.
  • S3-Rec (zhou2020s3) introduces self-supervised learning tasks to enrich representation learning, yet still relies on pure ID embeddings lacking transferability.
@@
  • VQ-Rec (hou2023learning) utilizes vector quantization for sequential recommendation.
  • MISSRec (wang2023missrec) aligns text and image features using pre-extracted embeddings to improve multimodal sequential recommendation.
@@
  • P5-CID (hua2023index) organizes multiple recommendation tasks in a text-to-text format and models different tasks uniformly using the T5 model.
  • VIP5 (geng2023vip5) extends P5 into a multimodal framework using CLIP image encoders and text prompts. These methods achieve strong zero-shot generalization but depend heavily on language priors instead of learned item semantics.
@@
  • TIGER (rajput2023recommender) is the first to employ RQ-VAE for constructing discrete semantic item IDs, enabling generative recommendation.
  • MQL4GRec (zhai2025multimodal) further quantizes multimodal embeddings into a unified “quantitative language” to enhance cross-domain transferability.

### v2

@@ 1. INTRODUCTION
Recommend systems have become indispensable across e-commerce, digital advertising, and content streaming, where they surface relevant items from massive catalogs and drive key business metrics such as CTR, retention, and revenue(Zhang et al., 2019a). Among various paradigms, sequential recommendation focuses on modeling users’ historical interactions to predict their next actions, thereby capturing evolving preferences and short-term intents. Early neural methods such as GRU4Rec(Jannach and Ludewig, 2017) leveraged recurrent architectures to encode temporal dynamics, while self-attention models like SASRec(Kang and McAuley, 2018) improved long-range dependency modeling and parallelism. Building on the Transformer, BERT4Rec(Sun et al., 2019) introduced bidirectional contextualization for sequences and set a strong foundation for many subsequent advances. These developments establish the Transformer as a de facto backbone for sequence modeling in recommendation.
@@
To enhance representation quality, recent work focuses on exploring multimodal signals—text (titles, descriptions), images, and structured attributes—to enrich item semantics beyond bare IDs (e.g., VBPR, MMGCN, DualGNN)(He and McAuley, 2016; Wei et al., 2019; Wang et al., 2021). In parallel, generative and prompt-based recommenders reframe recommendation as a language or sequence generation problem(Deldjoo et al., 2024), unifying diverse tasks through instruction tuning or tokenization (e.g., P5, M6-Rec, TIGER, VIP5)(Geng et al., 2022; Cui et al., 2022; Rajput et al., 2023; Geng et al., 2023). More recent work such as MQL4GRec(Zhai et al., 2025) and DiffMM(Jiang et al., 2024)further explored quantized or generative modeling to bridge continuous multimodal features and discrete semantics. MQL4GRec employs Residual Vector Quantization (RQ-VAE) (Zhai et al., 2025) to convert multimodal representations into discrete semantic tokens. However, it performs quantization independently for each modality, which results in inconsistent codebook distributions across modalities. This misalignment weakens the shared semantic space and hinders the integration of multimodal information in sequential modeling. DiffMM, while leveraging diffusion-based reconstruction(Ho et al., 2020) for generation, primarily focuses on denoising processes rather than fine-grained multimodal fusion. These examples reveal a fundamental challenge: existing models either perform static multimodal fusion without adaptivity, or decouple quantization from sequence modeling, thereby limiting both interpretability and generalization.
@@
The development of Residual Vector Quantization (RQ-VAE) techniques(van den Oord et al., 2017; Zeghidour et al., 2021) has further inspired us to address these limitations. Similarly to how pretraining–finetuning(Devlin et al., 2019) and prompt-tuning(Brown et al., 2020) revolutionized natural language generation by introducing a shared vocabulary with transferable semantics, we envision a comparable paradigm shift for recommendation. Instead of representing items by arbitrary, meaningless IDs, we propose to replace them with semantic IDs—compact sequences of quantized semantic tokens that encapsulate multimodal meaning. These semantic IDs act as both identifiers and language units, bridging the gap between discrete recommendation tokens and rich multimodal semantics. In this view, each token from the quantized codebook serves as a semantic atom, much like words in a vocabulary that encode transferable priors across domains. Consequently, recommendation can be reformulated as a language of items, where interactions are modeled as sentences composed of semantic tokens. This new representation not only enables cross-domain knowledge transfer and interpretability, but also maintains efficiency and scalability by being more compact than original modalities such as text or images. As illustrated in Figure 1, the proposed semantic-ID paradigm provides a unified, token-level bridge connecting multimodal content, quantized representations, and sequential reasoning.
@@
To this paper, we propose a novel Quantized BERT-style Multimodal Sequential Recommendation framework, named Q-Bert4Rec. Specifically, our method is motivated by the observation that traditional sequential recommenders represent items as discrete and meaningless IDs, which lack semantic understanding and hinder generalization across domains. To bridge this gap, Q-Bert4Rec introduces a unified three-stage framework that learns semantic IDs composed of quantized multimodal tokens, enabling both semantic richness and discrete modeling. In the first stage, we perform Cross-modal Semantic Injection, where a dynamic Transformer fuses textual, visual, and structural features into randomly initialized ID embeddings. A learnable gating mechanism adaptively controls the fusion depth, ensuring content-aware semantic enrichment for different items.In the second stage, we introduce Semantic Quantization using a Residual Vector Quantization module (RQ-VAE)(Zeghidour et al., 2021). The fused representations are discretized into compact, interpretable semantic tokens, which serve as the new identifiers—semantic IDs—to replace original item IDs. These quantized tokens form a unified vocabulary that encodes multimodal meaning in a transferable and efficient way. In the third stage, we design a Multi-mask Pretraining and Fine-tuning strategy to fully exploit sequential dependencies. By combining span, tail, and multi-region masking, the model captures both local continuity and long-range correlations in user behavior sequences. This stage enhances the model’s ability to reason over diverse sequential patterns and improves its adaptability to downstream recommendation tasks. We evaluate Q-Bert4Rec on multiple public Amazon benchmarks, and the results show significant improvements compared to strong baselines. The findings demonstrate that unifying semantic fusion and quantized modeling effectively bridges the gap between continuous multimodal representations and discrete recommendation reasoning.
@@ 2. RELATED WORK
Sequential recommendation. Sequential models aims to capture users’ dynamic interaction histories and predict their next potential actions. Early neural architectures, such as GRU4Rec (Jannach and Ludewig, 2017) and Caser (Tang and Wang, 2018), leveraged recurrent and convolutional structures to capture temporal patterns. Later, the emergence of self-attention mechanisms, particularly SASRec (Kang and McAuley, 2018), greatly improved long-range dependency modeling and parallel training efficiency. Building upon this, BERT4Rec (Sun et al., 2019) introduced a bidirectional Transformer encoder and masked item prediction, setting a strong foundation for subsequent Transformer-based recommenders.Recent works have explored different enhancements on this backbone. For instance, CL4SRec (Xie et al., 2022) and DuoRec (Qiu et al., 2022) introduced contrastive learning objectives to improve representation robustness, while FDSA (Zhang et al., 2019b) incorporated feature-level attention to exploit auxiliary attributes. Despite their success, these approaches still represent items using discrete, non-semantic IDs, which limits their transferability to unseen domains or new items. To address this, a new research trend focuses on generative recommendation, which reformulates recommendation as a language modeling or sequence generation problem. Representative works include P5 (Geng et al., 2022), M6-Rec (Cui et al., 2022), and TIGER (Rajput et al., 2023). These models leverage pre-trained language models (PLMs) and design prompt-based tasks to unify multiple recommendation objectives. P5 and M6-Rec transform recommendation into text generation through task prompts, while P5-ID (Hua et al., 2023) explore how the design of item identifiers affects model transferability. ColaRec (Wang et al., 2024) constructs generative item IDs by encoding collaborative signals, whereas TIGER(Rajput et al., 2023) is the first to introduce RQ-VAE(Zeghidour et al., 2021) for constructing discrete item IDs via vector quantization.
@@
Multimodal recommendation. Multi-modal models exploit side information—such as item text, images, and metadata—to provide richer representations of user–item interactions. Early works, such as VBPR (He and McAuley, 2016), enhanced matrix factorization with visual features extracted from product images. Subsequent research introduced graph-based models to capture modality relationships and high-order dependencies. For example, MMGCN (Wei et al., 2019) used graph neural networks to propagate multimodal signals; LATTICE (Zhang et al., 2021) built modality-specific item-item graphs for structural fusion; and DualGNN (Wang et al., 2021) modeled cross-modal attentions between user preferences and item content. Similarly, MVGAE (Yi and Chen, 2021) employed modality-specific variational autoencoders to integrate multimodal embeddings in a unified latent space. With the rapid advancement of large pre-trained models, multimodal recommendation has shifted toward leveraging foundation encoders for representation learning. VIP5 (Geng et al., 2023) extends the P5 framework by incorporating CLIP-based image encoders to align visual and textual modalities within a generative prompting framework. MMGRec (Liu et al., 2024) further utilizes a Graph RQ-VAE to construct item IDs from multimodal and collaborative signals, unifying item quantization and relational reasoning. IISAN (Fu et al., 2024) proposes a decoupled PEFT architecture for lightweight multimodal adaptation, demonstrating the scalability of pretrained multimodal encoders.
@@ 3.1. Problem Formulation
This bridges continuous multimodal semantics and discrete token-level modeling, enabling interpretability.
@@ 3.3. Dynamic Cross-Modal Semantic Injection
(5) $\mathcal{L}_{\text{align}}=\mathcal{L}_{it}+\mathcal{L}_{iv}+\mathcal{L}_{tv},$
where $\mathcal{L}_{it}$, $\mathcal{L}_{iv}$, and $\mathcal{L}_{tv}$ are symmetric InfoNCE losses between the fused, text, and image embeddings, The three terms encourage cross-modal alignment via contrastive learning
@@ 3.4. Semantic Quantization
Handling collisions. In practice, multiple items may be quantized to identical token sequences, leading to code collisions, where different items have the same token sequences. To mitigate this, we adopt the reallocation strategy proposed in MQL4GRec (Zhai et al., 2025). For $N$ colliding items, we first compute the distance tensor $\mathbf{D}\in\mathbb{R}^{N\times L\times K}$ between residual vectors and codewords in each level:
@@ 4.1. Experimental Setting
Dataset. We conduct experiments on the widely used Amazon Product Reviews dataset(Ni et al., 2019), which contains user–item interaction histories of different domains collected from Amazon over several years. Each domain represents a distinct product category and includes user reviews, ratings, and timestamps, making it suitable for evaluating cross-domain sequential recommendation.Statistics of these datasets are shown in Table  1.
@@
Evaluation Metrics. We evaluate model performance using two widely adopted ranking metrics: Hit Ratio (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K) with K = 1, 5, 10. Following previous work(Kang and McAuley, 2018), We use the leave-one-out for evaluation, where the most recent interaction is held out for testing and the second most recent for validation.
Implementation Details. We implement our model using PyTorch. For the dynamic cross-modal semantic injection module, we employ the pretrained LLaMA model to encode textual representations, and adopt the image encoder branch from CLIP (Radford et al., 2021), which uses ViT (Dosovitskiy, 2020) as its backbone, to extract visual features of items. The Transformer within this module is configured with a maximum depth of 5 layers.
@@
Following prior work (Sun et al., 2019), during pretraining, we apply three masking strategies with probabilities of 0.3 (span mask), 0.15 (multi-position mask), and 0.1 (tail mask). The model is trained with a batch size of 512 and a learning rate of 0.0005. The Transformer encoder consists of 4 layers with 4 attention heads. For fine-tuning, we use only the traditional mask strategy with a ratio of 0.3, a batch size of 256, and a learning rate of 0.0001. To prevent overfitting, we apply dropout with a rate of 0.2 to all Transformer layers.
@@ 4.2. Performance Comparison (RQ1)
Baselines. In this section, we compare our proposed approach with a wide range of sequential recommendation baselines, which are briefly described in Appendix A. The compared methods include GRU4Rec(Jannach and Ludewig, 2017), BERT4Rec (Sun et al., 2019), SASRec(Kang and McAuley, 2018), FDSA(Zhang et al., 2019b), S³-Rec(Zhou et al., 2020), VQ-Rec(Hou et al., 2023), MISSRec(Wang et al., 2023), P5-CID(Hua et al., 2023), VIP5(Geng et al., 2023), TIGER(Rajput et al., 2023), and MQL4Grec(Zhai et al., 2025). The experimental results are reported in Table 2.
@@
These consistent improvements validate that dynamic cross-modal semantic injection effectively captures richer multimodal dependencies, while the residual quantization module enhances semantic compactness and generalization across domains. Furthermore, the multi-mask pretraining strategy strengthens the model’s ability to understand diverse user interaction patterns, leading to better next-item prediction accuracy. Overall,in answer to RQ1, Q-BERT4Rec achieves the best performance in 13 out of 15 evaluation metrics, demonstrating strong generalization, and robustness across heterogeneous recommendation domains.
@@ Appendix A Baselines
  • GRU4Rec (Jannach and Ludewig, 2017) employs gated recurrent units to model sequential dependencies in user behaviors. Although effective for short sequences, it struggles to capture long-range dependencies and contextual semantics.
@@
  • BERT4Rec (Sun et al., 2019) applies bidirectional Transformers with the Masked Language Modeling (MLM) objective for next-item prediction. However, it models discrete item IDs without semantic meaning.
  • SASRec (Kang and McAuley, 2018) uses uni-directional self-attention for autoregressive sequence modeling. It effectively captures local dependencies but ignores item-level multimodal semantics.
  • FDSA (Zhang et al., 2019b) integrates user and item attributes into the attention mechanism to enhance personalization.
  • S3-Rec (Zhou et al., 2020) introduces self-supervised learning tasks to enrich representation learning, yet still relies on pure ID embeddings lacking transferability.
@@
  • VQ-Rec (Hou et al., 2023) utilizes vector quantization for sequential recommendation.
  • MISSRec (Wang et al., 2023) aligns text and image features using pre-extracted embeddings to improve multimodal sequential recommendation.
@@
  • P5-CID (Hua et al., 2023) organizes multiple recommendation tasks in a text-to-text format and models different tasks uniformly using the T5 model.
  • VIP5 (Geng et al., 2023) extends P5 into a multimodal framework using CLIP image encoders and text prompts. These methods achieve strong zero-shot generalization but depend heavily on language priors instead of learned item semantics.
@@
  • TIGER (Rajput et al., 2023) is the first to employ RQ-VAE for constructing discrete semantic item IDs, enabling generative recommendation.
  • MQL4GRec (Zhai et al., 2025) further quantizes multimodal embeddings into a unified “quantitative language” to enhance cross-domain transferability.